News
Apr
7
Emergency Maintenance – N4Cloud Portal – 09/04/17 20:00-23:59 GMT
Posted by Glenn Akester on 07 April 2017 12:36 PM

Background

On Sunday 7th April between 20:00 and 23:59 GMT Node4 engineers will be performing emergency maintenance on the firewalls which secure services accessed through the  N4Cloud Portal (portal.n4cloud.co.uk), to resolve a software issue.

 

Impact

N4Cloud portal access will be unavailable intermittently throughout the maintenance window.

Underlying cloud infrastructure and customer virtual machines will remain up and available throughout the maintenance.

 

Our Network Operations Engineers will closely monitor the work and will do everything possible to minimise any inconvenience to you. If you have any problems with N4Cloud portal access or functions after this time, or if you have any questions regarding the maintenance at any point, please call Technical Support at or visit www.node4.co.uk/support.

While it is important for Node4 to perform maintenance such as this to ensure the quality of our network, we do everything possible to minimise any inconvenience to our customers. We appreciate your patience and welcome any feedback.

 

Regards,

Node4 Technical Support


Read more »



Mar
30
Node4 3G/4G Network Issues – 30/03/17 13:30
Posted by Simon Bell on 30 March 2017 01:47 PM

We are currently investigating issue’s with our 3G/4G connectivity.

We Apologise for any inconvenience this is causing.

Regards

Node4 Technical Support

 

Update — 9:20am 31/03/2017:

The issue is only affect a few customer sites with SIM offline. We are still investigating this with 3G/4G service provider.

 

This issue was caused by EE’s outage and already resolved. If you still experiencing any problem, please contact Node4 Support on 0845 123 2229.


Read more »



Mar
30
London Global Switch 1 POP Planned Maintenance – 7/4/17 23:00-04:00
Posted by Simon Bell on 30 March 2017 01:41 PM

Background:

We intend to replace a core router with a new high powered router to increase our resiliency and capabilities at this POP

Services & Locations Affected:

London Global Switch 1 POP MPLS and N4-TRANSIT services

Impact:

POP is at risk for the duration of the maintenance, a small amount of customers will have direct downtime as a result of this maintenance, they will be contacted directly. We don’t anticipate any impact to customers outside of GS1. We will be re-routing traffic around this POP and moving traffic off Cogent London for a small period,
As such customers may notice reconvergance’s around this work including latency changes to certain destinations.

Maintenance Window:

Friday 7th April 2017 23:00 – 04:00

Preparation works will start 22:00, with service affecting work beginning at 23.00. We expect the work to take no longer than 2 hours however services will be “at risk” for the full maintenance window.

Maintenance Tracking will be available here

Apologies for any inconvenience this maintenance may cause.

Regards
Node4 Technical Support


Read more »



Mar
17
Emergency Maintenance – Node4 Derby Exchange – 17/03/2017
Posted by Zoe Yi Zhang on 17 March 2017 02:23 PM

17/03/2017 23:30 – 18/03/2017 00:00

Background:
Our Engineer will be performing scheduled maintenance on the Node4 Derby Exchange.

Impact
Customer’s conenctions terminated at Derby exchange will be affected due to the requirement of device reboot.
If you have any questions please contact Node4 Service Desk on 0845 123 2229 or www.node4.co.uk/support


Read more »



Mar
14
Partner Planned Maintenance – 23/03/17 23:00 GMT – NTT
Posted by Simon Bell on 14 March 2017 04:31 PM

Our provider NTT will be performing some short maintenance on our provided circuit.

We will take this opportunity to migrate the connection to a different router. Traffic will be gracefully routed around NTT whilst this takes place. Destinations currently reachable via NTT will move to Cogent and back again once the maintenance is complete. We do not expect any issue’s from this work, however transit via NTT will be at risk during the maintenance.

We expect the maintenance to last no longer than 1 hour with only a short duration.

We apologise for any inconvenience this may cause.

Update 22:55 – This maintenance is about to begin

Update 23:28 – Maintenance and migration of service has been completed. Traffic has been moved back onto NTT

Any issues from this work please raise with us so we can investigate immediately.

Regards

Node4 Technical Support


Read more »



Feb
15
Network Issues – 15/02/17
Posted by Alessandro Rizzo on 15 February 2017 09:47 PM

At 21:04 Node4 NOC has reported 10Gbps circuit down within our network, traffic has re-routed around resilient connections.

We are currently in talks with supplier to ascertain the reason for outage.

21:46 – Supplier is aware of the outage we are currently awaiting an update from the carrier.

23:54 – We have had reports from the supplier that there has been a fibre break. They are working to resolve this more updates follow.

01:10 – Our provider has updated us stated that engineers are due on site to the fibre break ETA 03:00am we have no ETA to resolve once they are on site more updates to follow.

03:40 – Our provider have advised that the engineers are now onsite at the fibre break and are working to fix this issue, further updates to follow.

05:20 – Engineers are currently still working to fix the fiber breaks, further updates to follow.

07:35 – Engineers are still working to fix this fiber break. Our provider is unable to provide an ETA for resolution at this time. Further updates to follow.

08:30 – Engineers are still working to fix this fiber break. No ETA as yet. We will update again in an hour.

09:30 – Provider incident Management team and techs are currently on a conference call, we expect an update imminently.

10:26 – Update from provider: – Engineers now need to gain access to a junction chamber located on the corner of a country road in order to continue working on this fiber break. This will require a traffic management team. Traffic management have been contacted and are on route. further updates to follow.

11:26 – Update from provider: –

All prep work is done, joint identified, specialist fibre engineers/OTDR’s etc on site,  still awaiting access to chamber via traffic management.

13:05 – Having chased our provider, we are informed that the status of this fault remains unchanged. Engineers on site are waiting on an emergency permit to work from the local authority. We are actively pursuing an ETA.

14:24 – Update from our providers: Permit and traffic will be in place from 15:00. Engineers are ready to work as soon as access is provided.

We will seek a further update and push for an ETA on fix at 17:00

17:30 – Our provider states that engineers are still on site, continuing to work to splice the fibres. Unfortunately no ETA is yet available on fix.

18:00 – Update from our provider :

The cable has multiple cuts beyond the chamber due to vandalism.

This means we will have to pull in approx. 650 metres of cable and re-splice both ends.

Current ETA for completion is 8 hours ETA approx. 02:00.

22:30 – Update from provider, Cable will be laid and prepped at 00:00. Splicing and testing will commence between 02:00 and 03:00.

02:20 – Update from provider, Cable preparation has been completed as scheduled. Splicing and testing is now taking place. Further updates to follow.

04:00 – Our provider has advised us that all 96 fibers have been successfully spliced, our monitoring is reflecting this as all circuits are up and passing traffic. Engineers are currently carrying our final checks and a further update will follow.

04:30 – Our provider has advised us that all testing has been completed and services have been resumed to normal operation. If you have any queries regarding this outage or have any outstanding issues please get in touch with our service desk at 08451232229.


Read more »